OpenStudy (anonymous):

Math Analysis: Prove the given identity sin 0/sin 0 + cos 0 = tan 0/1 + tan 0

OpenStudy (swag):

yes

OpenStudy (bahrom7893):

i think u're missing parenthesis

OpenStudy (anonymous):

i am going to guess that this says \[\frac{\sin(x)}{\sin(x)+\cos(x)}=\frac{\tan(x)}{1+\tan(x)}\]

OpenStudy (anonymous):

yeah! :)

OpenStudy (anonymous):

lol sorry ill write in (x) instead of (0) from now on.. xPP

OpenStudy (bahrom7893):

told u u were missing parenthesis... No not there, here: sin 0/(sin 0 + cos 0)

OpenStudy (anonymous):

that is my guess in any case and my second guess is that there is precious little trig here an amost all is algebra if you replace cosine by "a" and sine by "b" you get \[\frac{b}{a+b}=\frac{\frac{b}{a}}{1+\frac{b}{a}}\]

OpenStudy (anonymous):

multiply top and bottom of the left hand side by "a" and you get your identity in one step

OpenStudy (anonymous):

multiply top and bottom of the left hand side by "a" and you get your identity in one step

OpenStudy (anonymous):

sorry i meant the right hand side

OpenStudy (anonymous):

is b=1? or is that a=1?

OpenStudy (bahrom7893):

An easy way, just for the question posted: sin 0/(sin 0 + cos 0) = tan 0/(1 + tan 0) 0/(0+1) = 0/(1+0) 0/1 = 0/1 0=0 qed

OpenStudy (anonymous):

@bahrom7893 it is \[\theta\] not \[0\]

OpenStudy (bahrom7893):

OHHH LOL I was like.. hmm that's a joke hahahaha

OpenStudy (anonymous):

oh wait I get it tyvm <333!!!

OpenStudy (anonymous):

a is not 1, a is cosine b is sine

OpenStudy (anonymous):

i am a little slow

OpenStudy (anonymous):

point is that there is no trig here, it is just algebra

OpenStudy (anonymous):

no your fine, I was working on the equation while you were replying xD

OpenStudy (bahrom7893):

good point

OpenStudy (anonymous):

ty :)